OpenStudy (anonymous):

how to solve using synthetic division.... (2x^5+2x^4+4x^3-2x^3-4x+3)/x+2

OpenStudy (amistre64):

2x^2 right? thats a typo?

OpenStudy (amistre64):

(x+2); x = -2 -2| 2 +2 +4 -2 -4 +3 0 -4 +4 -16 +36 -64 --------------------- 2 -2 +8 -18 +32 -61

OpenStudy (anonymous):

yea thats suppose to be 4x^3-2x^2

OpenStudy (amistre64):

-61 2x^4 -2x^3 +8x^2 -18x +32 + ----- (x+2)

OpenStudy (anonymous):

amistre64 =right!

OpenStudy (amistre64):

yay!! .... i got one with out a typo lol

OpenStudy (amistre64):

do you see how I got it mcoley?

OpenStudy (anonymous):

lol wait okay im lost which is the answer lol im trying to follow i confused myself

OpenStudy (amistre64):

the first one is synthetic division to get all the "coefficiants" and a remainder: (x+2); x = -2 -2| 2 +2 +4 -2 -4 +3 0 -4 +4 -16 +36 -64 --------------------- 2 -2 +8 -18 +32 -61 Next we apply these coeefecients and the remainder into their normal form of: -61 2x^4 -2x^3 +8x^2 -18x +32 + ----- (x+2) It all depends on what format the "answer" is looking for really.

OpenStudy (amistre64):

i should first ask if you understand how we did the synthetic division... so do you?

OpenStudy (anonymous):

I see what your doing yea.. lol i get it

OpenStudy (amistre64):

good:)

OpenStudy (anonymous):

I was trying to do it myself n i must have miscalculated

OpenStudy (amistre64):

lol.... i miscalculate alot too :)